# Time-Dependent VAE for Building Latent Factor from Visual Neural Activity with Complex Dynamics

### v1

@@ 5 Discussion
This work presents a novel sequential VAE by introducing temporal dependence into a self-supervised generative model, aiming to reveal intrinsic correlations between neural activity from the visual cortex and visual stimulation during passive viewing. Our model, TiDe-SwapVAE, builds latent spaces conditioned on antecedent information to extract low-dimensional time-dependent representations from neural activity. Experiments on synthetic and mouse datasets demonstrate that our model constructs latent representations strongly correlated with the visual stimulation, in terms of decoding natural scenes or movie frames and extracting explicit temporal structure of neural dynamics.
Most latent variable models have focused on neural activity with task-driven behaviors. There is a paucity of research explaining the relationship between neural activity and visual stimuli with this type of model. [25] has made meaningful explorations of the response of the mouse visual cortex with CEBRA. Our work provides more comprehensive experiments of such data, illustrating the validity of latent variable models with temporal dependence for this area of research. However, neural responses elicited by complex natural visual stimuli show large variability across different subjects and different trials [33, 20]. Due to this, our model struggles to maintain stable performance to some extent. Improvements in model structures and learning algorithms are needed to mitigate this limitation.
Last but not least, our approach is not limited to mouse visual neural data and can be extended to neural data of other species or other modalities. As a promising computational model for neuroscience, it may also provide some insights to the field of machine learning.
Appendix A Backbone Structure of TiDe-SwapVAE
The encoder and decoder of our model are kept consistent with Swap-VAE. The encoder consists of three blocks, each of which is sequentially stacked with a linear layer, a batch normalization, and a ReLU activation. The output dimensions of the three linear layers is N, M, and M, where N is the number of neurons and M is the number of latent variables. The decoder is a symmetric structure of the encoder. We set the dimensions of the content and style spaces to be equal. For $\beta$-VAE, we use the same backbone structure. For other alternative models, we retain the structure of the original.
@@ Appendix A Backbone Structure of TiDe-SwapVAE
Appendix B Derivation of the Loss Function of TiDe-SwapVAE
Given that we use state variables and recurrent neural networks to build time dependence in VAE and process sequential data, we need to maximize the likelihood of the joint sequential distribution $p(\mathbf{x}_{1:T})$. Involving the latent variables $\mathbf{z}_{1:T}$, we have the variational lower bound:
@@ Appendix B Derivation of the Loss Function of TiDe-SwapVAE
Appendix C Characters of the Neural Dataset
In this work, we use a subset of the Allen Brain Observatory Visual Coding dataset [7, 27] recorded from six visual cortical regions of the mouse with Neuropixel probes. The full names and abbreviations of all cortical regions are listed in Table 3. We present the number of neurons for all chosen mouse.
Cortical Region Abbreviation Mouse 1 Mouse 2 Mouse 3 Mouse 4
primary visual cortex VISp 75 51 63 52
lateromedial area VISl 39 30 38 20
rostrolateral area VISrl 49 24 44 41
anterolateral area VISal 42 51 71 46
posteromedial area VISpm 62 90 19 64
anteromedial area VISam 94 72 60 64
@@ Appendix C Characters of the Neural Dataset
Appendix D Additional Visualization of Latent Representations
We visualize the latent representations of TiDe-SwapVAE in experiments of natural scenes for the other three mice and the latent representations of TiDe-SwapVAE, SwapVAE and CEBRA in experiments of natural movie stimuli for other parts of the movie. The results are in line with the conclusion in the main text.
(a) Mouse 2
(b) Mouse 3
(c) Mouse 4
(d) Mouse 2
(e) Mouse 3
(f) Mouse 4
Figure 5: Additional results on mouse visual cortex dataset under static natural scenes.
(a) 0-10s
(b) 0-10s
(c) 0-10s
(d) 20-30s
(e) 20-30s
(f) 20-30s
Figure 6: Additional results on mouse visual cortex dataset under natural movie.

### v2

@@ 5 Discussion
This work presents a novel sequential VAE by introducing time dependence into a self-supervised generative model, aiming to reveal intrinsic correlations between visual neural activity and visual stimulation. Our model, TiDeSPL-VAE, constructs latent representations conditioned on antecedent input to extract temporal relationships from neural activity in a natural way. Results of synthetic and mouse datasets demonstrate that our model outperforms alternative models and builds latent representations that are strongly correlated with visual stimulation, in terms of decoding natural scenes or movie frames and extracting explicit temporal structure from neural dynamics.
Most LVMs have focused on neural activity in the motor brain regions. There is a paucity of research explaining visual neural activity with this type of model. CEBRA has made meaningful explorations of neural responses of the mouse visual cortex [34]. Our work provides evidence that introducing time dependence with a chronological order plays a crucial role in studying visual neural activity with LVMs. However, there are some problems requiring further exploration. For example, neural responses elicited by complex natural visual stimuli show large variability across subjects and trials [43, 28]. Due to this, it is difficult for LVMs to consistently construct high-quality stimulus-correlated latent representations across conditions. Improvements in model structures and learning algorithms are needed in future work.
Last but not least, our approach is not limited to visual neural spikes from mice and can be extended to neural data from other brain regions and other species and of other modalities. As a promising model for neuroscience, it may also provide some insights into the field of machine learning.
Appendix A Backbone Structure of TiDeSPL-VAE
The encoder and decoder of our model are kept consistent with Swap-VAE. The encoder consists of three blocks, each of which is sequentially stacked with a linear layer, a batch normalization, and a ReLU activation. The output dimensions of the three linear layers are N, M, and M, where N is the number of neurons and M is the number of latent variables. The decoder is a symmetric structure of the encoder. We set the dimensions of the content and style spaces to be equal. For $\beta$-VAE, we use the same backbone structure. For other alternative models, we retain the structure of the original.
@@ Appendix A Backbone Structure of TiDeSPL-VAE
Appendix B Derivation of the Loss Function of TiDeSPL-VAE
Given that we use state factors and recurrent neural networks to build time dependence in VAE and process sequential data, we need to maximize the likelihood of the joint sequential distribution $p(\mathbf{x}_{1:T})$. Involving the latent variables $\mathbf{z}_{1:T}$, we have the variational lower bound:
@@ Appendix B Derivation of the Loss Function of TiDeSPL-VAE
Finally, we apply NT-Xent loss as the contrastive loss:
$\mathcal{L}_{\mathrm{contrast}}=-\log{\frac{\exp{\left(\mathrm{sim}\left(\mathbf{z}^{(s)},\mathbf{z}_{\mathrm{pos}}^{(s)}\right)/\tau\right)}}{\exp{\left(\mathrm{sim}\left(\mathbf{z}^{(s)},\mathbf{z}_{\mathrm{pos}}^{(s)}\right)/\tau\right)}+\sum{\exp{\left(\mathrm{sim}\left(\mathbf{z}^{(s)},\mathbf{z}_{\mathrm{neg}}^{(s)}\right)/\tau\right)}}}},$ (18)
where $\mathrm{sim}(*,*)$ is the cosine similarity and $\tau$ is the temperature coefficient.
Appendix C Synthetic Datasets
C.1 Generating Procedure
Non-temporal dataset
First, we generate labels $u_{i}$ from four uniform distributions on $[\frac{2i\times\pi}{4},\frac{(2i+1)\times\pi}{4}],i\in\{0,1,2,3\}$, in preparation for building four clusters. Second, for each cluster, we sample 2-dimensional latent variables $\mathbf{z}$ from independent Gaussian distribution with $(5\sin{u_{i}},5\cos{u_{i}})$ as mean and $(0.6-0.5|\cos{u_{i}}|,0.5|\cos{u_{i}}|)$ as variance. Third, we feed sampled latent variables into a RealNVP network [11] to form firing rates of 100-dimensional observations and generate the synthetic neural activity from the Poisson distribution. Each cluster contains 4,000 samples. All samples are split to 80% for training and 20% for test.
Temporal dataset
We generate three dynamic latent variables from the Lorenz system consisting of a set of nonlinear equations and the firing rates of 30 simulated neurons are computed by randomly weighted linear readouts from the Lorenz latent variables. The synthetic neural activity is also generated from the Poisson distribution. The hyperparameters of the Lorenz system follow some previous work [38]. We run the Lorenz system for 1s (1ms for a time point) from five randomly initialized conditions. Each condition contains 20 trials. We use 80% of the dataset (16 trials/condition) for training and 20% (4 trials/condition) for test.
C.2 Training Implementation
Non-temporal dataset
All models are set up with 32-dimensional latent variables and trained for 20,000 iterations with an optimizer of Adam and a learning rate of 0.0005. Notably, since this dataset does not involve time dependence, the length of input sequences is set to 1 even for models that can handle sequential data.
Temporal dataset
All models are set up with 8-dimensional latent variables and trained for 20,000 iterations with an optimizer of Adam and a learning rate of 0.001. Our model and LFADS use 50ms of data as input, while the other models take data at one time point because they can’t process sequential data.
Appendix D Characters of Neural Dataset
In this work, we use a subset of the Allen Brain Observatory Visual Coding dataset [9, 36] recorded from six visual cortical regions of the mouse with Neuropixel probes. The full names and abbreviations of all cortical regions are listed in Table 6. We present the number of neurons for all chosen mice.
@@ Appendix D Characters of Neural Dataset
Cortical Region Abbreviation Mouse 1 Mouse 2 Mouse 3 Mouse 4 Mouse 5
primary visual cortex VISp 75 51 93 63 52
lateromedial area VISl 39 30 56 38 20
rostrolateral area VISrl 49 24 58 44 41
anterolateral area VISal 42 51 43 71 46
posteromedial area VISpm 62 90 17 19 64
anteromedial area VISam 94 72 49 60 64
Appendix E Experiment Setup on Neural Dataset for Alternative Model
For $\beta$-VAE, pi-VAE, and Swap-VAE, each input sample is neural activity of an independent time point, both at the training and inference stages. For LFADS, we apply the same approach as our model. For CEBRA, following the original approach [34], we take the surrounding points centered on the target point to form a sequence of the same length as during the training stage (5 time points for natural scenes and 4 for natural movie) and compute its latent variables.
Appendix F Additional Results on Synthetic Non-Temporal Dataset
Figure 6: The regressed latent variables of alternatives on the synthetic non-temporal dataset.
Appendix G Additional Visualization of Latent Representations
We visualize latent representations of TiDeSPL-VAE in experiments under natural scenes for the other four mice and latent representations of TiDeSPL-VAE, SwapVAE and CEBRA in experiments under natural movie stimuli for other parts of the movie. The results are in line with the conclusion in the main text.
Figure 7: Visualization results of TiDeSPL-VAE on the visual neural dataset under natural scenes (Mouse 2-5).
Figure 8: Visualization results on the visual neural dataset under natural movie (Mouse 2). A-C. The 2-dimensional embedding of the latent representations for each frame in the first 10s of the movie. D-F. The 2-dimensional embedding of the latent representations for each frame in the last 10s of the movie. The first 10s show more entanglement across time, while the last 10s are rarely entangled. Nevertheless, our model still outperforms other alternative models.
Appendix H Additional Results of Ablation Studies
We present the results of ablation studies for the other three mice, which are consistent with the conclusion in the main text.
Table 7: The decoding scores of ablation studies on the loss function and the recurrent module of TiDeSPL-VAE. The standard error is computed based on 10 runs.
Natural Scenes Natural Movie
Models Mouse 3 Mouse 4 Mouse 5 Mouse 3 Mouse 4 Mouse 5
TiDeSPL-VAE 74.8$\pm$1.7 78.8$\pm$2.9 67.6$\pm$2.1 59.88$\pm$0.72 54.33$\pm$0.54 30.18$\pm$0.40
Without negative samples 67.2$\pm$2.5 76.0$\pm$2.0 59.6$\pm$2.6 45.67$\pm$0.60 44.17$\pm$0.42 19.93$\pm$0.35
Without contrastive loss 72.0$\pm$1.6 76.4$\pm$2.6 60.4$\pm$1.9 44.09$\pm$0.67 44.02$\pm$0.47 18.24$\pm$0.41
Without swap operation 66.4$\pm$1.9 72.4$\pm$2.4 65.2$\pm$3.0 45.30$\pm$0.41 43.12$\pm$0.57 21.83$\pm$0.39
Without swap operation and constrastive loss 59.6$\pm$2.7 66.8$\pm$2.3 53.6$\pm$2.8 22.33$\pm$0.81 26.49$\pm$0.66 12.02$\pm$0.49
With non-temporal prior 64.4$\pm$2.6 69.6$\pm$3.0 61.2$\pm$3.0 53.87$\pm$0.49 46.90$\pm$0.48 22.92$\pm$0.43
[…]
Table 8: The decoding scores (%) of ablation studies on the content and style latent representations of TiDeSPL-VAE. The standard error is computed based on 10 runs.
Natural Scenes Natural Movie
Mouse 3 Mouse 4 Mouse 5 Mouse 3 Mouse 4 Mouse 5
Content 76.4$\pm$1.6 77.6$\pm$2.3 68.0$\pm$2.3 65.92$\pm$0.51 57.07$\pm$0.44 32.39$\pm$0.48
Style 60.8$\pm$2.8 61.6$\pm$3.2 49.2$\pm$3.3 14.62$\pm$1.06 17.96$\pm$1.00 11.12$\pm$0.39
Figure 9: The results of ablation studies on the dimension of latent variables and the number of input neurons. Error bars indicate the standard error for 10 runs.